OpenStudy (anonymous):

If √6 is the geometric mean between 6 and another number, then the number is A.1 B.2 C.6

OpenStudy (anonymous):

G^2 = ab

OpenStudy (anonymous):

ie (sqrt (6) ) ^2 = 6 * b b = 1

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Welcome